Journalist Report – March 2th

Sol 18 – When space exploration brings philosophy to the table

“It is good to renew one’s wonder,” said the philosopher. “Space travel has again made children of us all.”

The Martian Chronicles by Ray Bradbury

“Alexandre to the Hab, I have arrived at the Observatory.”

When I heard this message over the radio this morning, I instinctively wrote it down in my notebook, for no particular reason other than that I was happy for Alexandre. The Sun was shining, he would be able to spend time in the solar observatory to record footage of our star and program his nighttime observations.

I also felt happy for Alice and Quentin: the Sun was shining, the hills were drying, boding well for tomorrow’s scheduled EVA. As the exploration EVA using the 3D map ended up being rescheduled to Sol 20 due to the weather, tomorrow, Quentin, Jérémy, and Alexandre will be looking for the checkpoints using only the 2D map.

I started preparing lunch while Jérémy and Corentin debriefed the Sociomapping reports received after the first half of the mission. This technique allows to literally “map out” the relationship between crewmembers: Using a series of weekly queries we answer regularly, a 3D map is generated, illustrating how the crew is organized, and how efficient communications are between us.

While Alexandre and Jérémy were preparing their notes, and memorizing paths and caches for tomorrow, Quentin made a discovery: the damaged weather station screen was in fact not needed! By placing a receiver near the Hab window, he was able to “hear” the weather station’s transmissions directly from inside the Hab and receive them on his computer.

Gathered around the Hab table for lunch, as we were rereading some questions sent by journalists about our mission and Martian exploration, we quickly slipped into debate mode: Why go to space? Why explore Mars? Many answers arose. Most of us mentioned science: good use of orbital space can help solve the climate crisis, studying Mars helps us understand our own Earth, even pure scientific curiosity can be a powerful motivator. But some crewmembers also mentioned a reason more difficult to justify, abstract yet almost visceral, even artistic: the desire to explore. Which is why I chose this quote, instead of going through Bradbury’s 18th chronicle as usual:

“‘It is good to renew one’s wonder,’ said the philosopher. ‘Space travel has again made children of us all.’”

Astronomy Report – March 2th

Astronomy Report

Name: Alexandre VINAS

Crew: 275
Date: 03/02/2023


Robotic Telescope Requested (choose one) MDRS-14 MDRS-WF

Objects Viewed: /

Problems Encountered: /


Solar Features Observed: Sun 230302 sunspots

Problems Encountered: My prominence capture was not very satisfying. The surrounding of the Sun was not black but very bright (Does it depend on the hour of the day ? I made this photo at about 12:00 pm). There was no prominence during my observation so that was not a problem as I did not use the prominence capture for the photo, but I would like to understand why, in order to fix the problem for the next times.

GreenHab Report – March 2th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 79.3°F

Average humidity: 26%

Hours of supplemental light: 0 hour

Daily water usage for crops: 10.30 gallons

Daily water usage for research and/or other purposes: 0 gallons

Water in the blue tank: 127.87 gallons

Time(s) of watering for crops: 9:35 & 17:35

Changes to crops: N/A

Narrative: Today was a pretty common day so I watered and took care of plants and crops.

However, as tomorrow is our media day, I made sure everything was into place, even though everything is always clean!

Apart from that, the aquaponics system is doing well, fish too, and all the plants as well.

Harvest: N/A

Support/supplies needed: N/A

Operations Report – March 2th

Crew 275 Operations Report 02Mar2023
SOL: 18

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9

Notes on non-nominal systems: None

Spirit rover used: No
• Hours: 209.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.5
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: No
• Hours: 221.7
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Nothing to report.

Summary of Hab operations: The Crew conducted routine operations. The lights of the kitchen hood were replaced. The light in the oven is still missing.
• WATER USE: 10.30
• Water: 418.17 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: Yes

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS. MDRScampus is used for experimental purpose.

Summary of suits and radios: Nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations. The new CO alarm was placed in the GreenHab.
• WATER USE: 10.30 gallons
• Heater: On
• Supplemental light: No
• Harvest: No

Summary of ScienceDome operations: The Crew Scientist conducted experiments on rock samples.
• Dual split: Heat / From 10 pm to 2 am

Summary of RAM operations: No operations were conducted in the RAM today. We retrieved the CO alarm. The

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support:
• There was a bigger leak in the roof of the RAM today, the window of the roof is slightly cracked, and drops of water came from this crack. I placed a bucket to collect the water underneath. The leak stopped in the evening.

Sol Summary Report – March 2th

Crew 275 Sol Summary Report 02Mar2023

Sol: 18

Summary Title: When space exploration brings philosophy to the table

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

With the precipitation of the last few days, we knew that the ground would be too muddy today to conduct EVAs, which is why we decided to wait a bit before performing our second EVA to Candor Chasma. It was apparently the right choice, as we can see from our windows that the mud is quickly drying out.

We conducted routine operations in the Hab, the GreenHab, and the Science Dome, finalizing most of the Human Factors experiments planned for Week 3. With the sunny weather, our Crew Astronomer was finally able to take his first pictures of the Sun. He has been waiting for this moment for so long and is very happy with the result!

The end of the afternoon was dedicated to cleaning the different modules, as we want to make a good impression on the BBC tomorrow. They have also sent us a list of questions that we have started to prepare for tomorrow’s interviews. Some of these questions were actually the topic of a very interesting discussion that we all had together after lunch.

Look Ahead Plan:

Tomorrow (Sol 19), we would like to have our second EVA to Candor Chasma. We then plan to go back there on Saturday (Sol 20). The EVA crews will have to find specific checkpoints based on a 2D map or a 3D render, respectively. We would like to take the opportunity of these EVAs to change the batteries of the weather instruments, since they cannot last for more than two days, especially in cold weather. In addition, since no precipitation is expected, we would like to re-install the LOAC next to the other atmospheric instruments.

The BBC is planned to arrive at the MDRS tomorrow (Sol 19) between 8:30 and 9:00. They will follow us on our EVA to Candor Chasma, then film inside the station in the afternoon, and leave at around 17:00.

Anomalies in work: None.

Weather: Sunny, low -1°C/30°F, high 7°C/44°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Request, Operations Report, Journalist Report, GreenHab Report, Astronomy Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site